OpenStudy (anonymous):

my last question to my test h.w given that f(x) = 2x^2 - 3x + 4 and g (x) = 3x^2 + 4x - 5, find: 1. (f+g)(-2) 2. (f-g)(-3) 3. (f . g) (2/3) 4. (f/g) (- 4/5) 5. (f o g) (3) 6. (gof)(-4) 7.(f o f) ( 3/2) 8. ( g o g) (-1/2)

OpenStudy (anonymous):

help functions discrete math

OpenStudy (amistre64):

what is "test hw"?

OpenStudy (anonymous):

home work test

OpenStudy (amistre64):

in what part are you stumped on?

OpenStudy (anonymous):

the prof did not teach this TT

OpenStudy (anonymous):

some of my classmate and friends confused to

OpenStudy (amistre64):

well; what does "(f+g)(-2)" mean? or rather, what do you understand this notation to mean?

OpenStudy (anonymous):

function + ....

OpenStudy (amistre64):

ok ... lets work with that the notation is a shorthand for: f(-2) + g(-2) ... but what do these notations signify?

OpenStudy (anonymous):

they are the same

OpenStudy (amistre64):

(f+g)(-2) means.. f(-2) + g(-2). but what does f(-2) mean? what does g(-2) mean? notice the 2 functions have been named and defined as: f(x) = 2x^2 - 3x + 4 and g(x) = 3x^2 + 4x - 5 f(-2) is a way to say; what does f(x) = when x=-2 g(-2) is a way to say: what does g(x) = when x=-2

OpenStudy (anonymous):

ic ic if im looking for X example sir (f+g) means x = (-2) (f-g) x = (-3) correct or wrong?

OpenStudy (amistre64):

if i read you correctly, then yes. whatever is in the ()'s is the value they want you to use for "x" in the functions

OpenStudy (amistre64):

the f+g part tells you want they want you to do with the functions; add them together in this case. (f-g) subtract them (f.g) = multiply them ..etc

OpenStudy (anonymous):

f(-2) 2-2^2 - 3^2+ 4 g(-2) 3-2^2 + 4-2-5 sir?

OpenStudy (amistre64):

the part that might be tricker is the (fog), (gog) and the like when there is the "o" in the middle

OpenStudy (amistre64):

youve got it :)

OpenStudy (anonymous):

after that sir what to do next?

OpenStudy (amistre64):

(f+g)(-2) = f(-2): 2(-2)^2 -3(-2) +4 (+) g(-2): 3(-2)^2 +4(-2) -5 just add up all your values now to find (f+g)(-2)

OpenStudy (anonymous):

sir 2(-2) =-4?

OpenStudy (amistre64):

(f+g)(-2) = f(-2): 2(-2)^2 -3(-2) +4 + g(-2): 3(-2)^2 +4(-2) -5 ------------------------ 5(4) +1(-2) -1 = 20 -2-1 = 17 if i see it right

OpenStudy (amistre64):

2(-2) = -4 yes

OpenStudy (amistre64):

but; 2(-2)^2 = 2(4) = 8

OpenStudy (anonymous):

i understand sir very well everything u tell your good 2x to my prof

OpenStudy (amistre64):

thanx :) the notation of (fog)(x) has a special meaning it means that we use one function as the x value for the other function ... whcih can be tricky

OpenStudy (amistre64):

we see that f(x) where x = -2 means f(-2) but: f(g(x)) means that for every value of x in f(x) we use the value of g(x) at x=-2. does that make sense?

OpenStudy (anonymous):

i copy these before u go i have my test paper here need to send tommorow before prelim exam hahaha

OpenStudy (anonymous):

sir 5(4) + 1 = 21 right then - 2 = 19? or wrong?

OpenStudy (amistre64):

f(g(-2)) means we find g(-2) like normal: g(-2) = 3(-2)^2 + 4(-2) - 5 = 3(4) + 4(-2) - 5 = 12 - 8 - 5 = 4 - 5 = -1 g(-2) = -1 f(g(-2)) means f(-1) f(g(-2)) = 2(g(-2))^2 - 3(g(-2)) + 4 = 2(-1)^2 - 3(-1) + 4 = 2(1) + 3 + 4 = 2 + 7 = 9 all of these mean the same thing ... (fog)(-2) = 9 f(g(x)) at x=-2; = 9 f(g(-2)) = 9 f(-1) = 9

OpenStudy (amistre64):

5(4) + 1 20 + 1 21 ....... is correct

OpenStudy (anonymous):

so no.1 question answer is 18 right or wrong?

OpenStudy (amistre64):

but ,, that is not what its asking for :) 5 times (4) plus, 1 times (-2), minus 1 is how we read it 5 times (4) = 5(4) = 20 20 plus, 1 times (-2) = 20 + (-2) = 18 18 minus 1 = 18-1 = 17

OpenStudy (anonymous):

so i get -1 from the last +4 -5 =-1 sir correct or wrong?